OpenStudy (anonymous):

What are the solutions of 4x2 = -20x? Answer x = 0, x = 5 x = 0, x = -5 x = 4, x = 5 x = 4, x = -5

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4x^2 + 20x = 0 }\) Factorise. \( \color{Black}{\Rightarrow 4x(x + 5) = 0 }\) Using the zero product rule, either \(4x = 0\), or \(x + 5 = 0\). Simplifying, \(x = 0\) or \(-5\)

OpenStudy (anonymous):

thank you

Parth (parthkohli):

yw